# Reasoning or Simply Next Token Prediction? A Benchmark for Stress-Testing Large Language Models

### v1

@@ 4.1 Evaluation Protocol
We evaluated three large language models: gpt-3.5-turbo, gemini-1.0-pro, and llama3-8b. The assessments for gpt-3.5-turbo and gemini-1.0-pro were conducted using the Gemini-benchmark pipeline [1]. For these models, we set the temperature parameter to 0 and utilized carefully crafted prompts that required responses in the format of “Answer: Letter of Choice.” This approach ensures that the generated responses are directly comparable and suitable for evaluation. Additionally, both models were evaluated in the 5-shot setting, using examples from our development dataset to enhance their contextual understanding. llama3-8b was evaluated using the lm-evaluation-harness framework [8]. This model employed a different evaluation strategy; it uses log likelihood to determine the model’s responses. Consistent with the other models, llama3-8b also uses the same 5-shot setting, ensuring a standardized comparison across all tests.
@@ 4.2 Results and Analysis
General Trend. Table 1 shows the accuracy of the three models gpt-3.5-turbo, gemini-1.0-pro, and llama3-8b evaluated in both MMLU and our MMLU-SR. The data highlights how each model performs in the Humanities, Social Sciences, STEM, and Other academic fields, providing average scores for each subset. We observe consistent drop in model performance across all subsets when transitioning from the standard MMLU dataset to the more challenging MMLU-SR dataset, as evidenced by the decline in average accuracy from 0.677 on the MMLU dataset to 0.616, 0.520, and 0.459, on our MMLU-SR’s “Question Only”, “Answer Only”, and “Question and Answer” subsets, respectively, for the gpt-3.5-turbo model. This trend of decreased performance is similarly observed in the other models.
Table 1: Performance of gpt-3.5-turbo, gemini-1.0-pro, and llama3-8b.
Model Humanities Social Sciences STEM Other Average
GPT-3.5-turbo
MMLU (5-shot) 0.723 0.770 0.554 0.714 0.677
Question Only (5-shot) 0.661 0.702 0.506 0.641 0.616
Answer Only (5-shot) 0.540 0.595 0.441 0.538 0.520
Question and Answer (5-shot) 0.469 0.523 0.396 0.476 0.459
Gemini-1.0-pro
MMLU (5-shot) 0.728 0.758 0.596 0.703 0.686
Question Only (5-shot) 0.687 0.744 0.539 0.658 0.645
Answer Only (5-shot) 0.619 0.670 0.504 0.591 0.586
Question and Answer (5-shot) 0.582 0.622 0.472 0.544 0.546
Llama3-8B
MMLU (5-shot) 0.593 0.757 0.557 0.729 0.651
Question Only (5-shot) 0.546 0.685 0.507 0.668 0.595
Answer Only (5-shot) 0.455 0.599 0.460 0.557 0.510
Question and Answer (5-shot) 0.421 0.538 0.424 0.499 0.465
Model Humanities Social Sciences STEM Other Average
GPT-3.5-turbo
Question Only (5-shot) 8.58% 8.83% 8.67% 10.22% 9.08%
Answer Only (5-shot) 25.31% 22.73% 20.40% 24.65% 23.27%
Question and Answer (5-shot) 35.12% 32.08% 28.52% 33.30% 32.26%
Gemini-1.0-pro
Question Only (5-shot) 5.63% 1.85% 9.56% 6.40% 5.86%
Answer Only (5-shot) 14.96% 11.61% 15.44% 15.91% 14.48%
Question and Answer (5-shot) 20.05% 17.94% 20.81% 22.60% 20.85%
Llama3-8B
Question Only (5-shot) 7.92% 9.51% 8.98% 8.36% 8.69%
Answer Only (5-shot) 23.27% 20.87% 17.41% 23.56% 21.28%
Question and Answer (5-shot) 28.16% 28.93% 23.88% 31.56% 28.63%
@@
Accuracy Drop in Each Category. Table 2 shows several aspects in the relative percentage drop of accuracy in MMLU-SR compared to that in MMLU across different categories for gpt-3.5-turbo, gemini-1.0-pro, and llama3-8b:
@@
    Humanities and Social Sciences. For gpt-3.5-turbo, the accuracy drops significantly in the Humanities category, with a slightly lower drop in Social Sciences. The gemini-1.0-pro model demonstrates notable resilience in the Social Science category within the “Question Only” dataset, exhibiting only a minimal performance decrease. In contrast, there are moderate drops in MMLU-SR’s “Answer Only” and “Question and Answer” datasets. Despite these decreases, gemini-1.0-pro still shows the smallest performance decline in the Humanities and Social Science categories compared to the other two models evaluated. llama3-8b exhibits a similar pattern to gpt-3.5-turbo, with the Humanities and Social Sciences categories showing a moderate percentage drop, though slightly lower than gpt-3.5-turbo in the “Question and Answer” dataset.
@@
    STEM. For both gemini-1.0-pro and llama3-8b, the STEM category shows the least decrease in accuracy across the MMLU-SR datasets. This indicates that these models maintain their performance better in STEM subjects compared to other categories when faced with symbol replacement. However, gpt-3.5-turbo does not follow this trend and instead shows a more significant drop in the STEM category, especially in the “Answer Only” and “Question and Answer” datasets. This suggests that gpt-3.5-turbo may be less robust and more sensitive to modifications in the STEM domain.
@@
    Other. The Other category generally shows a significant drop across all models and datasets, with the highest drops often observed in the “Question and Answer” dataset. This suggests that questions in this category might be more context-dependent and thus more affected by symbol replacement.
Detailed Accuracy Drop in Each Subject. Table 3 shows a detailed comparison of accuracy scores across different models evaluated on various subjects in the Humanities category. The MMLU scores serve as a baseline for comparison. Generally, gemini-1.0-pro performs best across most subjects in this category, closely followed by gpt-3.5-turbo and then llama3-8b. All models achieve relatively high accuracy scores in Humanities, with performance generally around 0.6 to 0.7 for most subjects. However, all models seem to struggle with solving problems in Moral Scenarios, where the accuracy score is strikingly lower, around 0.2 to 0.3, very close to the random baseline of 0.25.
Subject MMLU Question Only Answer Only Question and Answer GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 Formal Logic 0.492 0.437 0.476 0.413 0.373 0.413 0.381 0.413 0.413 0.310 0.381 0.310 Logical Fallacies 0.767 0.816 0.730 0.724 0.761 0.712 0.650 0.736 0.601 0.589 0.663 0.472 Moral Disputes 0.769 0.740 0.740 0.688 0.697 0.645 0.569 0.645 0.584 0.442 0.604 0.552 Moral Scenarios 0.342 0.278 0.403 0.280 0.400 0.336 0.245 0.210 0.246 0.255 0.242 0.244 Philosophy 0.768 0.759 0.736 0.666 0.656 0.601 0.450 0.521 0.524 0.357 0.492 0.441 World Religions 0.860 0.848 0.836 0.819 0.801 0.819 0.643 0.690 0.667 0.608 0.655 0.608 High School European History 0.800 0.806 0.782 0.764 0.788 0.764 0.612 0.739 0.667 0.545 0.709 0.636 High School Us History 0.858 0.848 0.848 0.833 0.799 0.760 0.730 0.730 0.637 0.623 0.711 0.588 High School World History 0.844 0.882 0.827 0.802 0.852 0.772 0.620 0.768 0.667 0.557 0.747 0.612 Prehistory 0.738 0.815 0.722 0.682 0.747 0.679 0.549 0.664 0.586 0.469 0.614 0.528 International Law 0.843 0.835 0.851 0.777 0.802 0.826 0.628 0.802 0.736 0.595 0.769 0.653 Jurisprudence 0.806 0.843 0.769 0.676 0.731 0.731 0.556 0.667 0.593 0.417 0.556 0.417 Professional Law 0.512 0.557 0.460 0.471 0.528 0.449 0.391 0.464 0.360 0.334 0.430 0.358
@@
Subject MMLU Question Only Answer Only Question and Answer GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 Econometrics 0.535 0.579 0.500 0.465 0.482 0.430 0.395 0.439 0.342 0.395 0.386 0.316 High School Macroeconomics 0.651 0.738 0.636 0.608 0.654 0.592 0.464 0.569 0.503 0.379 0.477 0.433 High School Microeconomics 0.769 0.815 0.723 0.660 0.739 0.597 0.508 0.668 0.542 0.429 0.559 0.462 High School Government And Politics 0.912 0.943 0.891 0.845 0.881 0.850 0.674 0.751 0.710 0.591 0.720 0.642 Public Relations 0.700 0.718 0.709 0.664 0.645 0.664 0.618 0.600 0.636 0.555 0.536 0.536 Security Studies 0.698 0.808 0.735 0.645 0.780 0.678 0.539 0.678 0.624 0.510 0.657 0.600 Us Foreign Policy 0.860 0.890 0.870 0.780 0.840 0.860 0.600 0.720 0.720 0.610 0.720 0.690 Human Sexuality 0.817 0.260 0.779 0.763 0.740 0.695 0.649 0.687 0.649 0.542 0.687 0.611 Sociology 0.846 0.866 0.866 0.796 0.851 0.816 0.746 0.821 0.806 0.632 0.766 0.706 High School Geography 0.864 0.859 0.813 0.763 0.813 0.763 0.697 0.712 0.596 0.581 0.662 0.576 High School Psychology 0.862 0.879 0.831 0.802 0.820 0.771 0.705 0.760 0.670 0.602 0.706 0.596 Professional Psychology 0.722 0.737 0.725 0.637 0.683 0.608 0.546 0.634 0.518 0.456 0.592 0.459
@@
Table 5 shows a detailed comparison of accuracy across various STEM subjects for different models. Each model demonstrated varying degrees of success across the subjects, with notable difficulties in some areas. Particularly, the subject with the lowest accuracies among all models is College Mathematics, where the accuracy is consistently around 0.3 in MMLU-SR’s “Question and Answer” dataset, reflecting the challenges in solving math problems. High School Mathematics also shows similar trends, highlighting the models’ struggles with subjects involving extensive calculations and problem-solving.
Subject MMLU Question Only Answer Only Question and Answer GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 Abstract Algebra 0.420 0.360 0.310 0.320 0.340 0.300 0.380 0.340 0.310 0.360 0.390 0.320 College Mathematics 0.250 0.430 0.350 0.260 0.350 0.350 0.180 0.340 0.380 0.270 0.340 0.340 High School Statistics 0.537 0.625 0.537 0.421 0.537 0.477 0.394 0.481 0.440 0.352 0.435 0.417 Elementary Mathematics 0.484 0.463 0.423 0.452 0.410 0.368 0.463 0.460 0.392 0.431 0.392 0.347 High School Mathematics 0.289 0.426 0.404 0.300 0.381 0.359 0.344 0.381 0.415 0.319 0.344 0.370 Astronomy 0.770 0.822 0.697 0.671 0.770 0.625 0.579 0.645 0.579 0.500 0.645 0.539 College Physics 0.461 0.412 0.461 0.373 0.412 0.333 0.373 0.333 0.373 0.324 0.324 0.333 Conceptual Physics 0.694 0.672 0.579 0.532 0.600 0.506 0.464 0.532 0.383 0.319 0.426 0.332 High School Physics 0.377 0.464 0.444 0.391 0.371 0.384 0.258 0.331 0.305 0.291 0.298 0.291 College Chemistry 0.470 0.470 0.460 0.450 0.490 0.400 0.400 0.420 0.400 0.410 0.370 0.380 High School Chemistry 0.527 0.537 0.527 0.473 0.512 0.502 0.409 0.517 0.463 0.325 0.429 0.429 College Biology 0.792 0.854 0.771 0.715 0.778 0.736 0.542 0.701 0.625 0.403 0.653 0.562 High School Biology 0.813 0.845 0.771 0.755 0.823 0.710 0.606 0.700 0.639 0.542 0.703 0.577 College Computer Science 0.560 0.590 0.490 0.510 0.490 0.460 0.450 0.530 0.340 0.470 0.450 0.400 Computer Security 0.800 0.790 0.800 0.770 0.760 0.780 0.550 0.660 0.680 0.520 0.710 0.650 High School Computer Science 0.710 0.740 0.680 0.730 0.710 0.620 0.660 0.700 0.610 0.600 0.650 0.580 Machine Learning 0.411 0.509 0.527 0.411 0.384 0.527 0.438 0.473 0.402 0.312 0.482 0.330 Electrical Engineering 0.614 0.717 0.648 0.579 0.586 0.628 0.448 0.531 0.483 0.379 0.462 0.483
[…]
Subject MMLU Question Only Answer Only Question and Answer GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 GPT Gemini Llama3 Anatomy 0.674 0.667 0.711 0.556 0.644 0.615 0.519 0.504 0.467 0.481 0.437 0.430 Clinical Knowledge 0.758 0.762 0.766 0.694 0.706 0.649 0.608 0.615 0.574 0.543 0.551 0.521 College Medicine 0.699 0.665 0.642 0.624 0.688 0.566 0.532 0.572 0.590 0.497 0.572 0.486 Human Aging 0.726 0.753 0.717 0.583 0.632 0.610 0.457 0.614 0.462 0.363 0.561 0.336 Medical Genetics 0.790 0.790 0.780 0.680 0.730 0.730 0.580 0.630 0.640 0.520 0.580 0.570 Nutrition 0.765 0.761 0.765 0.693 0.729 0.712 0.516 0.641 0.631 0.428 0.598 0.582 Professional Medicine 0.776 0.713 0.717 0.706 0.647 0.643 0.551 0.603 0.585 0.467 0.540 0.507 Virology 0.524 0.548 0.560 0.428 0.464 0.476 0.398 0.518 0.452 0.349 0.500 0.380 Business Ethics 0.740 0.680 0.610 0.610 0.640 0.620 0.540 0.540 0.520 0.470 0.530 0.490 Management 0.816 0.835 0.883 0.796 0.786 0.845 0.650 0.728 0.718 0.583 0.641 0.631 Marketing 0.910 0.889 0.893 0.872 0.855 0.833 0.722 0.739 0.645 0.654 0.714 0.598 Global Facts 0.440 0.420 0.350 0.390 0.380 0.360 0.380 0.460 0.350 0.290 0.310 0.370 Miscellaneous 0.884 0.835 0.837 0.840 0.803 0.775 0.709 0.670 0.599 0.614 0.644 0.539 Professional Accounting 0.489 0.521 0.496 0.496 0.504 0.486 0.365 0.440 0.362 0.397 0.440 0.369
@@
5 Discussion
5.1 Limitation
While MMLU-SR provides a novel approach to evaluate language models by incorporating symbol replacement, it has two major limitations: Firstly, some questions and answers remain unchanged without symbol replacement, especially in cases involving straightforward mathematical numbers or simple question forms. While these unchanged elements should not introduce significant noise, their presence might affect the intended complexity and variability of the evaluation. Secondly, MMLU-SR only involves multiple-choice questions, limiting its utility in assessing models’ capabilities in generative tasks. This restriction could hinder comprehensive evaluations of a model’s ability to generate coherent and contextually appropriate responses beyond structured choices. As the field progresses, expanding MMLU-SR to include generative components could provide a more holistic assessment of language models.
5.2 Future Work
For future work, it would be interesting to enhance MMLU-SR by creating more structured and domain-specific evaluations. One approach is to develop a pilot exam-style dataset, where a set of questions refers to a single chart or diagram, allowing us to study the model’s understanding within specific contexts more deeply. Additionally, it would also be interesting to introduce a “Hard” version of MMLU-SR by replacing terms with their synonyms or antonyms as dummy words. An example is shown in Table 12 from Appendix D. This will challenge the models’ domain understanding and ability to discern subtle differences in meaning, thereby providing a more rigorous evaluation of their reasoning capabilities. Furthermore, one could also modify our development dataset for each subset by integrating CoT prompts, enabling MMLU-SR to test whether 5-shot CoT improves the evaluation scores in larger-scale models, rather than simply performing zero-shot CoT using ChatGPT-4.
5.3 Conclusion
@@ Appendix C System Instruction Example
Model Humanities Social Sciences STEM Other Average

### v2

@@ 4.1 Evaluation Protocol
We evaluated seven models across OpenAI, Gemini, Llama families: gpt-3.5-turbo, gpt-4o-mini, gpt-4o, gemini-1.0-pro, gemini-1.5-pro, llama3-8b, and llama3-70b. The evaluation for GPT and Gemini models was conducted using the Gemini-benchmark pipeline Akter et al. (2023). For these models, we set the temperature parameter to 0 and utilized carefully crafted prompts that required responses in the format of “Answer: Letter of Choice.” This approach ensures that the generated responses are directly comparable and suitable for evaluation. Additionally, both models were evaluated in the 5-shot setting, using examples from our development dataset to enhance their contextual understanding. Llama3 was evaluated using the lm-evaluation-harness framework Gao et al. (2023). This model employed a different evaluation strategy; it uses log likelihood to determine the model’s responses. Consistent with the other models, Llama3 also uses the same 5-shot setting, ensuring a standardized comparison across all tests. The complete results of all seven models are available in Appendix E.
@@ 4.2 Results and Analysis
General Trend. Table 1 shows the accuracy of the four models gpt-4o-mini, gpt-4o, gemini-1.5-pro, and llama3-70b evaluated in both MMLU and our MMLU-SR. The data highlights how each model performs in the Humanities, Social Sciences, STEM, and Other academic fields, providing average scores for each subset. We observe consistent drop in model performance across all subsets when transitioning from the standard MMLU dataset to the more challenging MMLU-SR dataset, as evidenced by the decline in average accuracy from 0.771 on the MMLU dataset to 0.710, 0.655, and 0.585, on our MMLU-SR’s “Question Only”, “Answer Only”, and “Question and Answer” subsets, respectively, for the gpt-4o-mini model. This trend of decreased performance is similarly observed in the other models.
@@
Accuracy Drop in Each Category. Table 2 shows several aspects in the relative percentage drop of accuracy in MMLU-SR compared to that in MMLU across different categories for gpt-4o-mini, gpt-4o, gemini-1.5-pro, and llama3-70b:
@@
    Humanities and Social Sciences. For gpt-4o-mini and gpt-4o, the accuracy drops significantly in the Humanities category, with a slightly lower drop in Social Sciences. The gemini-1.5-pro shows the smallest performance decline in the Humanities and Social Science categories compared to the other two models evaluated. llama3-70b exhibits a pattern similar to gpt-4o-mini, with the Humanities and Social Sciences categories showing a moderate percentage drop, though slightly higher than gpt-4o-mini, in the “Answer Only” and “Question and Answer” dataset.
@@
    STEM. For gemini-1.5-pro and llama3-70b, the STEM category shows a relatively moderate decrease in accuracy across the MMLU-SR datasets. Notably, gemini-1.5-pro experiences the highest drop of 12.72% in the “Question Only” dataset, indicating some sensitivity in this area. llama3-70b demonstrates a similar trend, with the highest drop of 9.47% in the STEM category, suggesting both models retain some robustness in STEM but are still impacted by symbol replacement. On the other hand, gpt-4o-mini experiences a higher drop in the “Answer Only” and “Question and Answer” datasets, particularly with a 22.93% drop in the latter, highlighting its relative vulnerability in this domain compared to gemini-1.5-pro and llama3-70b.
@@
    Other. The Other category generally shows a significant drop across all models and datasets, with the highest drops often observed in the “Question and Answer” dataset. For example, gpt-4o-mini experiences a notable drop of 25.19%, the highest among all categories and models, indicating a high sensitivity to contextual changes in this area. Similarly, llama3-70b follows closely with a 27.40% drop, which is the highest in the Other category for this model. gemini-1.5-pro also shows a substantial drop of 16.44%, though slightly less compared to the other models, suggesting that the “Other” category, like Humanities, might be more context-dependent and hence more susceptible to performance degradation when symbols are replaced.
Detailed Accuracy Drop in Each Subject. Table 3 shows a detailed comparison of accuracy scores across different models evaluated on various subjects in the Humanities category. The MMLU scores serve as a baseline for comparison. gpt-4o demonstrates exceptional performance across most subjects in this category, often leading in accuracy, particularly in complex subjects like Philosophy and International Law. gemini-1.5-pro also shows strong performance, but gpt-4o frequently matches or exceeds its accuracy. Notably, gpt-4o performs particularly well in subjects like High School World History and Jurisprudence. However, all models continue to struggle with Moral Scenarios, where the accuracy score drops significantly, particularly for llama3-70b, which shows a drastic decrease, reflecting a higher sensitivity to the challenges posed by the MMLU-SR datasets
@@
Table 5 shows a detailed comparison of accuracy across various STEM subjects for different models. Each model demonstrated varying degrees of success across the subjects, with notable difficulties in some areas. College Mathematics and High School Mathematics remain challenging for all models, including gpt-4o, with accuracy dropping to around 0.4 to 0.5 in MMLU-SR’s “Question and Answer” dataset. However, gpt-4o shows marked improvement in subjects like Astronomy, College Biology, and High School Biology, maintaining high accuracy even in the more challenging MMLU-SR datasets. The subject with the lowest accuracies among all models is still High School Mathematics, where llama3-70b struggles the most, especially in the Answer Only” and “Question and Answer” datasets. Similarly, College Physics and Abstract Algebra also show significant drops in accuracy across all models, highlighting the persistent challenges in subjects involving extensive calculations and complex problem-solving.
Table 6 shows a detailed comparison of accuracy scores across different models evaluated on various subjects in the Other category. We observe that gpt-4o performs exceptionally well in MMLU, with accuracy consistently above 0.9 in most subjects, significantly outperforming other models. Marketing stands out with a particularly high accuracy of 0.949 for gpt-4o, indicating outstanding performance in this subject. Professional Accounting shows improved performance with gpt-4o, achieving an accuracy of 0.766 in MMLU. Virology remains challenging, but gpt-4o shows improvement with an accuracy of 0.578. While there is still a drop in accuracy from MMLU to MMLU-SR’s “Question and Answer” dataset, gpt-4o maintains relatively high performance, with accuracy generally staying above 0.7 for most subjects. Even in challenging areas like Virology and Global Facts, gpt-4o demonstrates resilience, maintaining accuracy levels significantly higher than other models.
@@
5 Conclusion
@@
6 Acknowledgements
We thank the anonymous reviewers for their thoughtful feedback and for recognizing the significance and contributions of our research. This research is based on work partially supported by DARPA Grant HR00112420363, NSF Grant IIS-2127918, and NSF CAREER Award IIS-2340125.
@@ Appendix C System Instruction Example
Dataset Humanities Social Sciences STEM Other Average
@@ Appendix D Hard MMLU-SR Example
Table 13: Complete performance of gpt-3.5-turbo, gpt-4o-mini, gpt-4o, gemini-1.0-pro, gemini-1.5-pro, llama3-8b, and llama3-70b.
Dataset Humanities Social Sciences STEM Other Average
GPT-3.5-turbo
MMLU (5-shot) 0.723 0.770 0.554 0.714 0.677
Question Only (5-shot) 0.661 0.702 0.506 0.641 0.616
Answer Only (5-shot) 0.540 0.595 0.441 0.538 0.520
Question and Answer (5-shot) 0.469 0.523 0.396 0.476 0.459
GPT-4o-mini
MMLU (5-shot) 0.793 0.858 0.689 0.782 0.771
Question Only (5-shot) 0.744 0.792 0.621 0.724 0.710
Answer Only (5-shot) 0.659 0.738 0.602 0.651 0.655
Question and Answer (5-shot) 0.588 0.666 0.531 0.585 0.585
GPT-4o
MMLU (5-shot) 0.880 0.906 0.771 0.854 0.845
Question Only (5-shot) 0.838 0.856 0.702 0.811 0.792
Answer Only (5-shot) 0.764 0.824 0.705 0.760 0.757
Question and Answer (5-shot) 0.708 0.754 0.635 0.712 0.695
Gemini-1.0-pro
MMLU (5-shot) 0.728 0.758 0.596 0.703 0.686
Question Only (5-shot) 0.687 0.744 0.539 0.658 0.645
Answer Only (5-shot) 0.619 0.670 0.504 0.591 0.586
Question and Answer (5-shot) 0.582 0.622 0.472 0.544 0.546
Gemini-1.5-pro
MMLU (5-shot) 0.849 0.881 0.802 0.815 0.832
Question Only (5-shot) 0.795 0.836 0.700 0.754 0.764
Answer Only (5-shot) 0.741 0.816 0.747 0.739 0.758
Question and Answer (5-shot) 0.690 0.752 0.670 0.681 0.694
Llama3-8B
MMLU (5-shot) 0.593 0.757 0.557 0.729 0.651
Question Only (5-shot) 0.546 0.685 0.507 0.668 0.595
Answer Only (5-shot) 0.455 0.599 0.460 0.557 0.510
Question and Answer (5-shot) 0.421 0.538 0.424 0.499 0.465
Llama3-70B
MMLU (5-shot) 0.681 0.868 0.697 0.814 0.765
Question Only (5-shot) 0.635 0.812 0.631 0.770 0.712
Answer Only (5-shot) 0.539 0.683 0.565 0.622 0.602
Question and Answer (5-shot) 0.523 0.653 0.536 0.591 0.576
Table 14: Complete relative percentage drop of accuracy in MMLU-SR compared to MMLU.
Dataset Humanities Social Sciences STEM Other Average
GPT-3.5-turbo
Question Only (5-shot) 8.58% 8.83% 8.67% 10.22% 9.08%
Answer Only (5-shot) 25.31% 22.73% 20.40% 24.65% 23.27%
Question and Answer (5-shot) 35.12% 32.08% 28.52% 33.30% 32.26%
GPT-4o-mini
Question Only (5-shot) 6.18% 7.69% 9.87% 7.42% 7.91%
Answer Only (5-shot) 16.90% 13.99% 12.63% 16.75% 15.05%
Question and Answer (5-shot) 25.85% 22.38% 22.93% 25.19% 24.12%
GPT-4o
Question Only (5-shot) 4.77% 5.52% 8.95% 5.03% 6.27%
Answer Only (5-shot) 13.18% 9.05% 8.56% 11.01% 10.41%
Question and Answer (5-shot) 19.55% 16.78% 17.64% 16.63% 17.75%
Gemini-1.0-pro
Question Only (5-shot) 5.63% 1.85% 9.56% 6.40% 5.86%
Answer Only (5-shot) 14.96% 11.61% 15.44% 15.91% 14.48%
Question and Answer (5-shot) 20.05% 17.94% 20.81% 22.60% 20.85%
Gemini-1.5-pro
Question Only (5-shot) 6.36% 5.11% 12.72% 7.48% 8.17%
Answer Only (5-shot) 12.72% 7.38% 6.86% 9.33% 8.89%
Question and Answer (5-shot) 18.73% 14.64% 16.46% 16.44% 16.59%
Llama3-8B
Question Only (5-shot) 7.92% 9.51% 8.98% 8.36% 8.69%
Answer Only (5-shot) 23.27% 20.87% 17.41% 23.56% 21.28%
Question and Answer (5-shot) 28.16% 28.93% 23.88% 31.56% 28.63%
Llama3-70B
Question Only (5-shot) 6.75% 6.45% 9.47% 5.41% 6.93%
Answer Only (5-shot) 20.85% 21.31% 18.94% 23.59% 21.31%
Question and Answer (5-shot) 23.20% 24.77% 23.10% 27.40% 24.71%
Appendix E Complete Experiment Results
Table 13 shows our complete experiment results of different LLMs including gpt-3.5-turbo, gpt-4o-mini, gemini-1.0-pro, gemini-1.5-pro, llama3-8b and llama3-70b. The percentage drop of each model is shown in Table 14.
Appendix F Numbers of Human Modified Terms
Figure 5 shows the number of manually modified term definitions across 41 subject glossaries. These glossaries were created by consolidating related topics from the original 57 subjects in the MMLU dataset. Of the 28,676 terms initially generated by gpt-3.5-turbo for these 41 subjects, approximately 1,197 (4.2%) required human modification.
Figure 5: Comparison of total generated terms (red) and human-modified terms (blue) across 41 subject glossaries